4. What is the boiling point of a solution that contains 12.5g biphenyl, C12H10, in 100g of bromobenzene? The normal boiling point of bromobenzene is 156°C, and kb for bromobenzene is +6.26°C/m.
1
Expert's answer
2021-10-22T03:53:13-0400

"ΔT = k_bm \\
k_b = +6.26 \;°C/m \\
M(biphenyl) = 12 \times 12 + 10 =154 \;g/mol \\
n(biphenyl) = \frac{12.5}{154}=0.0811 \;mol \\
Molality \;m= \frac{0.0811}{0.1} = 0.811 \;mol/kg \\
ΔT = 6.26 \times 0.811 = 5.08 \;°C \\
T = 156 + 5.08 = 161.08 \;°C"